# ETF: An Entity Tracing Framework for Hallucination Detection in Code Summaries

### v3

@@ Appendix D String Matching Heuristics
The above regexes are designed to capture all the cases of entity forms in summary. It can be noted that these regex rules are evaluated in a single module.
@@ E.2 Guidelines
The annotation process was conducted in two stages. In the first stage, we implemented a three-step procedure to annotate hallucinations in code summaries independent of specific hallucination categories.
The first part of the annotation process involved validating the Named Entity Recognition Output. This involved annotating missed or incorrectly identified entities in the summary itself by selecting the appropriate label from the drop-down.
The second part of the annotation involved evaluating if each sentence accurately describes the entity in the code snippet, marking the entity-sentence pair as:
@@
The third part involved rating the summary based on hallucination severity as
@@
The second stage involved defining hallucination categories based on annotator feedback and organizing them into a structured taxonomy (Figure 2). This finalized taxonomy was then provided to the annotators, who were asked to assign a specific hallucination category from the predefined options. Annotators were strongly encouraged to include comments explaining their annotations, as these explanations can be useful for researchers utilizing our dataset.

### v4

@@ D.1 String Matching Heuristics
The above regexes are designed to capture all the cases of entity forms in summary, evaluated in a single module.
D.2 Prompt based Approaches
Here, we discuss the general prompt-based approach we tried for Intent detection. We give the complete prompt in Figure 8. Qualitatively, we observed the following drawbacks of this approach:
  • We observe high Hallucination in the generated output, which leads to the introduction of fabricated sentences not present in the summaries.
  • We observed inaccurate extraction of the sentences, where the extracted sentence has slight variations from the original sentence present in the summary.
  • We observe missing sentences, i.e., not all the sentences discussed in the summary are captured in the generated output.
Intent Detection Prompt Assume you are a Java expert. You have to identify all the relevant sentences about the given entity. Here, a sentence is relevant to the mapped entity if the sentence discusses the given entity. You have to generate the output strictly in the JSON format: $\{{entity\_name:``",relevant\_sentence:``"}\}$ Given Entity: {mapped entity} Given Summary: {SUMMARY}
Figure 8: Intent Detection Prompt- This prompt was used to retrieve all the relevant sentences from the summary
These observations led us to prefer simpler string-matching heuristics, which are significantly cheaper in computational aspects.
@@ E.2 Guidelines
The annotation process was conducted in two stages. In the first stage, we implemented a three-step procedure to annotate hallucinations in code summaries independent of specific hallucination categories. The first step of the annotation process involved validating the Named Entity Recognition Output. This involved annotating missed or incorrectly identified entities in the summary itself by selecting the appropriate label from the drop-down.
The second step of the annotation involved evaluating if each sentence accurately describes the entity in the code snippet, marking the entity-sentence pair as:
@@
The third step involved rating the summary based on hallucination severity as
@@
The second stage involved defining hallucination categories based on annotator feedback and organising them into a structured taxonomy (Figure 2). This finalised taxonomy was then provided to the annotators, who were asked to assign a specific hallucination category from the predefined options. Annotators were also encouraged to include comments explaining their annotations, as these explanations can be useful for researchers utilising our dataset.